OpenStudy (anonymous):

Multiply and simplify. 2/3*(1/3+3 1/4) A.4 7/18 B.4 7/9 C.2 7/9 D.2 7/18

OpenStudy (tkhunny):

What's preventing you from following the instructions? The mixed number?

OpenStudy (agl202):

2/3 × (1/3 + 3 1/4) = 2/3 (1/3 + 13/4) = 2/3 (43/12) = 43/18 = 2 7/18